OpenStudy (ebydaqueen19):

SOMEONE PLZ HELP!!!! CALCULATE THE MASS(KG) OF WATER PRODUCED FROM THE COMBUSTION OF 1.0 GALLON (3.8L) OF GASOLINE(C8H18). THE DENSITY OF GAS IN .79 G/ML.

OpenStudy (anonymous):

ok start by writing combustion reaction, then we will continue...

OpenStudy (ebydaqueen19):

ok i thin kit is C8H18(l) +O2(g) = H2O(l) ?

OpenStudy (ebydaqueen19):

did i get the states of the compounds right?

OpenStudy (anonymous):

nope thats not right... C8H18 + O2 -> CO2 + H2O now is that equation ok like that (dont get in to states...)

OpenStudy (ebydaqueen19):

ok thank you so now whats next

OpenStudy (paxpolaris):

well we don't care how much CO2 is being produced ...it is just important to note all the H in C8H18 goes into H2O.

OpenStudy (paxpolaris):

use unit analysis to get the answer....

OpenStudy (ebydaqueen19):

ok so now in order to figure out the mass of the water produced what information would i look at

OpenStudy (paxpolaris):

you start with 3.8 L C8H8 ... convert that into kg

OpenStudy (ebydaqueen19):

then to kg of water?

OpenStudy (ebydaqueen19):

or do u multiply kg of gas by density of gas?

OpenStudy (paxpolaris):

\[3.8 l\ C_8H_{18} \times {.79 kg\ C_8H_{18} \over l\ C_8H_{18}}\] kg of C8H18

OpenStudy (anonymous):

density of gas = mass of gas / volume of gas

OpenStudy (paxpolaris):

then you want to get how much of that mass is Hydrogen

OpenStudy (ebydaqueen19):

oh then from mass of H to mass of H2O

OpenStudy (ebydaqueen19):

and also kryten wrote a chem equation above can you check that to see if its correct?

OpenStudy (ebydaqueen19):

cuz he said you dnt have to worry about the states of the compounds in the reaction...is tht only for precipitation reactions you have to worry about tht?

OpenStudy (anonymous):

you dont have to worry about states because you dont have to calculate anything state related example combustion energy or enthalpy etc.

OpenStudy (paxpolaris):

\[3.8 \cancel{l\ C_8H_{18}} \times {.79 kg\ C_8H_{18} \over \cancel{l\ C_8H_{18}}} \times {18\ kg\ H \over114\ kg\ C_8H_{18}}\] ... 114 is the Molecular mass of C8H18 out of which 18 comes from H

OpenStudy (paxpolaris):

this gives you the mass of hydrogen in kg that you are using to make the water ... you just need to convert it into the mass of water

OpenStudy (ebydaqueen19):

so you would divide by atomic mass for water?

OpenStudy (paxpolaris):

\[3.8 \cancel{l\ C_8H_{18}} \times {.79\ \cancel{kg\ C_8H_{18}} \over \cancel{l\ C_8H_{18}}} \times {18\ kg\ H \over114\ \cancel{kg\ C_8H_{18}}} \times {18\ kg\ H_2O \over 2\ kg\ H}\]

OpenStudy (paxpolaris):

...the atomic mass of water divided by the amount that comes from water. and there's your answer. :)

OpenStudy (ebydaqueen19):

how did u get 18kg h2o

OpenStudy (ebydaqueen19):

i mean how did u know

OpenStudy (anonymous):

molar mass

OpenStudy (paxpolaris):

Hydrogern mass = 1 Oxygen mass = 16 H2O = 18 .... out of which 2 comes from hydrogen

OpenStudy (paxpolaris):

for density of C8H18 i used kg/L which is the same as g/ML. multiply all those numbers and i got ~4.3 kg of water

OpenStudy (ebydaqueen19):

THANK YOU SO MUCH FOR ALL YOUR HELP